# VecFusion: Vector Font Generation with Diffusion

### v1

@@ 1 Introduction
††Equal contributionVector fonts are extensively used in graphic design, arts, publishing, and motion graphics. As opposed to rasterized fonts, vector fonts can be rendered at any resolution without quality degradation, and can be edited via intuitive control point manipulation. However, authoring high-quality vector fonts remains a challenging and labor-intensive task, even for expert designers. Recent approaches [27, 6, 46] use VAEs or autoregressive models to automatically synthesize vector fonts, but they often struggle to capture a diverse range of topological structures and glyph variations, due to the inherent ambiguity of vector curves. As a result, they frequently create artifacts and imprecise control point positions, compromising the overall quality and editability of the synthesized fonts.
@@ 3.1 Raster diffusion stage
To encode the font style, we experimented with two approaches depending on the application. To generate missing glyphs in incomplete fonts, we create a one-hot vector representation for all font styles available in our dataset. Given a target font style, its one-hot vector is mapped to a continuous embedding through a learned look-up table as above. To generate glyphs conditioned on a few exemplar images, we concatenate the input images channel-wise and pass them through a convnet to get a font style feature map (see Appendix section: A: Implementation Details).
@@
The denoiser network predicts the per-channel noise component of the input image, which is also a -channel image (see Appendix section: A: Implementation Details).
@@
In all our experiments, we use the following hyperparameters. Following [32], we set the number of diffusion steps to 1000 and used cosine noise schedule in the forward diffusion process. We used the AdamW optimizer [28] with learning rate . The feature embeddings for character identifiers are set to be -dimensional. The control points are rendered as Gaussian blobs with radius of pixels. The raster image resolution is set to . Lower resolutions cause increasing overlaps between the rendered blobs, making the control point field more ambiguous. Increasing the resolution increases the computational overhead for the raster denoiser. The above resolution represented a good trade-off, as we practically found in our experiments. As mentioned above, we use channels to encode control point ordering and multiplicity as colors. We practically observed that channels were enough to guide the vector diffusion stage. Depending on the dataset, fewer channels could be used instead e.g., in cases of glyphs with few control points, or no multiplicities. In general, these hyperparameters can be adjusted for different vector graphics tasks – our main point is that the raster images and fields are useful as guidance to produce high-fidelity vector fonts, as shown in our ablation.
@@ 3.2 Vector diffusion stage
The denoiser for this stage is formulated as an encoder-only transformer [10], which takes the tensor as input and is conditioned on the embedding of time step , and the generated raster image from the raster diffusion model. We use a ConvNet to encode the raster image into high-dimensional features, which are input to the transformer via cross-attention similar to [37]. The transformer predicts the noise content as a tensor at each step A: Implementation Details.
@@
At test time, given a sampled tensor from unit Gaussian noise and a generated raster image of the previous stage, the denoiser network is applied in a series of steps to generate the target tensor . Following the Analog Bits approach [8], the discrete binary bits in the target tensor representation are modeled as real numbers. These are simply thresholded to obtain the final binary bits at inference time. Given the predicted path membership, we create a set of vector paths according to the largest generated control path index number. Each non-null entry in the generated tensor yields a control point. The control points are implicitly ordered based on their entry index. The location of the control point is defined as the coordinate center of the assigned cell in the generated tensor plus the predicted relative displacement. Given this generated information, we directly reconstruct the vector paths without requiring any further refinement or post-processing. 3
@@
In our implementation, we set the upper bound for the number of control points to , which was sufficient for the datasets we experimented with. We use bits to represent the path membership, which can support up to distinct vector paths. This was also sufficient for the datasets in our experiments. We set to , resulting in grid cells which can be represented by binary bits. Together with the two-dimensional relative displacement, the final dimension of our target tensor is in our experiments. Similar to our raster diffusion model, we set the number of diffusion steps to 1000, used cosine noise schedule, and the AdamW optimizer [28] with learning rate . During testing, we use the DDPM sampler [15] with steps.
@@ 4.1 Generation of missing Unicode glyphs
Table 1 shows the quantitative results for ChiroDiff and the alternative variants of our method on the full test set. The full version of our method outperforms ChiroDiff and our reduced variants on all evaluation metrics. We note that a glyph in one family variation e.g., italics might have a different variation in the same font, e.g., bold. Although two different family variations of the same glyph often have largely different control point locations and distributions, we create an even more challenging subset of the test set where we removed all test glyphs that had a different family variation in the same font within the training set. The results on this subset are reported in 2. Chirodiff and other variants still have much higher error than ours. This indicates that our two-stage approach, and the mixed discrete-continuous representation of vector paths along with the control point fields are all important to achieve high performance.
| Method | CD | #cp diff | #vp diff | |
| Ours | 0.069 | 0.46 | 15.05 | 0.033 |
| DeepVecFont-v2 [46] | 0.098 | 1.05 | 34.84 | 0.052 |
| Input reference | (a) | (b) | (c) | |||
@@
Figure 6 shows qualitative comparisons for missing glyph generation. Compared to our method, we observe that ChiroDiff produces imprecise control points and curve structure, resulting in significant distortions and artifacts in the synthesized glyphs. We also observed degraded results in all alternative variants of our method in terms of misplaced or skipped control points. Figures 1,5 show additional results of missing glyph generation for our method on the Google Font dataset for various target fonts and characters. Figure 7 demonstrates multiple samples generated by our diffusion pipeline with random seeds. The samples adhere to the same font style, while having subtle variation in the glyph proportions and control point distributions. From a practical point of view, a designer can explore multiple such samples, and choose the most preferred variant.
In Figure 12, instead of using the vector diffusion model, we use off-the-shelf vectorization methods on the raster glyph image produced by the raster diffusion stage. As shown in the comparison, this approach often fails to produce coherent curve topology and structure.
@@ 4.2 Few-shot font style transfer
For this application, we compare with DeepVecFont-v2 [46], which previously demonstrated few-shot font style transfer. To perform the comparison, we use the dataset proposed in the DeepVecFont-v2 paper. The dataset includes lowercase and uppercase Latin characters in various font styles – there are total training fonts and test ones. Each test case contains reference characters from a novel font (unobserved during training). The reference characters are available in both vector and raster format. Methods are supposed to transfer this novel font style to testing characters. We note that DeepVecFont-v2 requires the vector representation of the reference characters as additional input condition, while our method only uses the raster reference images. From this aspect, our method can be considered as disadvantaged in this comparison.
@@
Table 3 shows numerical comparisons based on the same evaluation metrics as in the missing glyph generation application. Despite the disadvantage, our method outperforms DeepVecFont-v2 on all metrics.
@@
Figure 8 demonstrates font style transfer results. We observed that DeepVecFont-v2 tends to succeed in capturing the font style of the reference glyphs, yet still often produces subtle distortions in the vector paths. Our method produces results that match the style of references glyphs with less artifacts, even in challenging styles. Additional results are in Figure 1 and 11.
@@ 5 Conclusion
We require vector paths as supervision and cannot take advantage of raster images as additional supervision that may be available. Another limitation is shown in Figure 10. Cinzel is an uppercase font. For the missing glyph ₫(dong), a lowercase glyph, our method produces an uppercase glyph. While the generated glyph preserves the style of the font and the structure of the glyph (the stroke), it does not preserve the symbolic meaning of the glyph.
@@
Generative models of images in the raster domain have been successfully used to learn complex priors about the visual world that can be used in various tasks such as inpainting and 3D tasks. We believe that a generative model for vector fonts could also have additional downstream applications, such as automatic completion of vector logos or icons, or be extended to produce 3D parametric curves or surfaces.

### v2

@@ 1 Introduction
†† ∗Equal contributionProject website: https://vikastmz.github.io/VecFusion/
Vector fonts are extensively used in graphic design, arts, publishing, and motion graphics. As opposed to rasterized fonts, vector fonts can be rendered at any resolution without quality degradation, and can be edited via intuitive control point manipulation. However, authoring high-quality vector fonts remains a challenging and labor-intensive task, even for expert designers. Recent approaches [31, 6, 51] use VAEs or autoregressive models to automatically synthesize vector fonts, but they often struggle to capture a diverse range of topological structures and glyph variations, due to the inherent ambiguity of vector curves. As a result, they frequently create artifacts and imprecise control point positions, compromising the overall quality and editability of the synthesized fonts.
@@ 3.1 Raster diffusion stage
To encode the font style, we experimented with two approaches depending on the application. To generate missing glyphs in incomplete fonts, we create a one-hot vector representation for all font styles available in our dataset. Given a target font style, its one-hot vector is mapped to a continuous embedding through a learned look-up table as above. To generate glyphs conditioned on a few exemplar images, we concatenate the input images channel-wise and pass them through a convnet to get a font style feature map (see supplement for more details.)
@@
The denoiser network predicts the per-channel noise component of the input image, which is also a -channel image (see supplement for more details.).
@@
In all our experiments, we use the following hyperparameters. Following [36], we set the number of diffusion steps to 1000 and used cosine noise schedule in the forward diffusion process. Training takes days on 8 A100 GPUs. We used the AdamW optimizer [32] with learning rate . The feature embeddings for character identifiers are set to be -dimensional. The control points are rendered as Gaussian blobs with radius of pixels. The raster image resolution is set to . Lower resolutions cause increasing overlaps between the rendered blobs, making the control point field more ambiguous. Increasing the resolution increases the computational overhead for the raster denoiser. The above resolution represented a good trade-off, as we practically found in our experiments. As mentioned above, we use channels to encode control point ordering and multiplicity as colors. We practically observed that channels were enough to guide the vector diffusion stage. Depending on the dataset, fewer channels could be used instead e.g., in cases of glyphs with few control points, or no multiplicities. In general, these hyperparameters can be adjusted for different vector graphics tasks – our main point is that the raster images and fields are useful as guidance to produce high-fidelity vector fonts, as shown in our ablation.
@@ 3.2 Vector diffusion stage
The denoiser for this stage is formulated as an encoder-only transformer [10], which takes the tensor as input and is conditioned on the embedding of time step , and the generated raster image from the raster diffusion model. We use a ConvNet to encode the raster image into high-dimensional features, which are input to the transformer via cross-attention similar to [41]. The transformer predicts the noise content as a tensor at each step.
@@
At test time, given a sampled tensor from unit Gaussian noise and a generated raster image of the previous stage, the denoiser network is applied in a series of steps to generate the target tensor . Following the Analog Bits approach [8], the discrete binary bits in the target tensor representation are modeled as real numbers. These are simply thresholded to obtain the final binary bits at inference time. Given the predicted path membership, we create a set of vector paths according to the largest generated control path index number. Each non-null entry in the generated tensor yields a control point. The control points are implicitly ordered based on their entry index. The location of the control point is defined as the coordinate center of the assigned cell in the generated tensor plus the predicted relative displacement. Given this generated information, we directly reconstruct the vector paths without requiring any further refinement or post-processing.
@@
In our implementation, we set the upper bound for the number of control points to , which was sufficient for the datasets we experimented with. We use bits to represent the path membership, which can support up to distinct vector paths. This was also sufficient for the datasets in our experiments. We set to , resulting in grid cells which can be represented by binary bits. Together with the two-dimensional relative displacement, the final dimension of our target tensor is in our experiments. Similar to our raster diffusion model, we set the number of diffusion steps to 1000, used cosine noise schedule, and the AdamW optimizer [32] with learning rate . Training is done separately from the raster diffusion stage and takes days on 8 A100 GPUs. During testing, we use the DDPM sampler [15] with steps. Generating a glyph by executing both stages takes around 10 seconds on a A100 GPU.
@@ 4.1 Generation of missing Unicode glyphs
Table 1 shows the quantitative results for ChiroDiff and the alternative variants of our method on the full test set. The full version of our method outperforms ChiroDiff and our reduced variants on all evaluation metrics. We note that a glyph in one family variation e.g., italics might have a different variation in the same font, e.g., bold. Although two different family variations of the same glyph often have largely different control point locations and distributions, we create an even more challenging subset of the test set where we removed all test glyphs that had a different family variation in the same font within the training set. The results on this subset are reported in Table 2. Chirodiff and other variants still have much higher error than ours. This indicates that our two-stage approach, and the mixed discrete-continuous representation of vector paths along with the control point fields are all important to achieve high performance.
Input reference | (a) | (b) | (c) | (d) | |||
@@
Figure 6 shows qualitative comparisons for missing glyph generation. Compared to our method, we observe that ChiroDiff produces imprecise control points and curve structure, resulting in significant distortions and artifacts in the synthesized glyphs. We also observed degraded results in all alternative variants of our method in terms of misplaced or skipped control points. Additional results are also in the supplementary. Figures 1,5 show additional results of missing glyph generation for our method on the Google Font dataset for various target fonts and characters. Figure 7 demonstrates multiple samples generated by our diffusion pipeline with random seeds. The samples adhere to the same font style, while having subtle variation in the glyph proportions and control point distributions. From a practical point of view, a designer can explore multiple such samples, and choose the most preferred variant.
In the supplementary, instead of using the vector diffusion model, we use off-the-shelf vectorization methods on the raster glyph image produced by the raster diffusion stage. As shown in the comparison, this approach often fails to produce coherent curve topology and structure.
@@ 4.2 Few-shot font style transfer
For this application, we compare with DeepVecFont-v2 [51] and DualVector [30], both of which previously demonstrated few-shot font style transfer. To perform the comparison, we use the dataset proposed in the DeepVecFont-v2 paper. The dataset includes lowercase and uppercase Latin characters in various font styles – there are total training fonts and test ones. Each test case contains reference characters from a novel font (unobserved during training). The reference characters are available in both vector and raster format. Methods are supposed to transfer this novel font style to testing characters. We note that DeepVecFont-v2 requires the vector representation of the reference characters as additional input condition, while DualVector and our method only use the raster reference images.
@@
Table 3 shows numerical comparisons based on the same evaluation metrics as in the missing glyph generation application. Our method outperforms DeepVecFont-v2 and DualVector on all metrics.
@@
Figure 8 demonstrates font style transfer results. We observed that both DeepVecFont-v2 tends to succeed in capturing the font style of the reference glyphs, yet still often produces subtle distortions in the vector paths. Our method produces results that match the style of references glyphs with less artifacts, even in challenging styles. Additional results are in Figure 1 and in the supplementary.
@@ 5 Conclusion
We require vector paths as supervision and cannot take advantage of raster images as additional supervision that may be available. Another limitation is shown in Figure 10. Cinzel is an uppercase font. For the missing glyph ₫(dong), a lowercase glyph, our method “hallucinates” an uppercase glyph. While the generated glyph preserves the style of the font and the structure of the glyph (the stroke), it does not preserve the symbolic meaning of the glyph.
@@
Generative models of images in the raster domain have been successfully used to learn complex priors about the visual world that can be used in various tasks such as inpainting and 3D tasks. We believe that a generative model for vector graphics in general could have several downstream applications, such as automatic generation and completion of logos, icons, line drawings and illustrations [20, 28, 29], or be extended to produce 3D parametric curves and surfaces [43].
Acknowledgments.
Our project was funded by Adobe Research.
@@ Acknowledgments.
Computation cost.
Raster-DM and Vector-DM are trained separately. Each of them is trained on 8 A100 GPUs for 5 days. Finally, at inference time, generating a glyph takes around 10 seconds on a A100.
@@
Additional comparisons with DeepVecFont-v2 [51].
Additional comparisons with ChiroDiff [9].
Please see Figure 13 for more comparisons with ChiroDiff [9] on the task of missing unicode generation.
Input reference | GT | NNs | Ours | DVF-v2 | |